OpenStudy (anonymous):

giving medal! plz help.. Rewrite the expression using exponents. Then evaluate. -6*-6+3*3*3-7 A. (-6)^2+(3)^2-7 ;14 B. -6(2)+3(3)-7 ; 13 C. (-6)^2+(3)^3-7 ; 56 D.-6(2)+3(3)-7; -10

OpenStudy (anonymous):

-6*-6 =(-6)^2 3*3*3=(3)^3 So, -6*-6+3*3*3-7 =(-6)^2+(3)^3-7

OpenStudy (anonymous):

(-6)^2=36 (3)^3=27 So the expression is = 36+27-7=56

OpenStudy (anonymous):

thank you